### v1

# Calibration Attention: Instance-wise Temperature Scaling for Vision Transformers

Abstract
Probability calibration is critical when Vision Transformers (ViTs) are deployed in risk-sensitive applications. However, the conventional solution—post-hoc temperature scaling—employs a single global scalar and requires tuning on a separate validation set. We propose Calibration Attention (CalAttn), a novel drop-in module designed to dynamically learn an adaptive, per-instance temperature directly from the ViT’s CLS token. Empirical evaluations on CIFAR-10/100, MNIST, TinyImageNet, ImageNet1K demonstrate that CalAttn substantially reduces calibration error—by up to a factor of 4—in models such as ViT-224, DeiT, and Swin Transformers, while introducing fewer than 0.1% additional parameters. Notably, the learned temperatures exhibit tight clustering around 1.0, starkly contrasting the large global values typically needed by standard temperature scaling methods. Calibration Attention offers a straightforward, efficient, and architecture-agnostic strategy for producing trustworthy and well-calibrated probabilities in state-of-the-art Vision Transformers.
1 Introduction
A classifier is well-calibrated when its predicted probabilities reflect the true likelihood of correctness: among all predictions made with 80% confidence, roughly 80% should in fact be correct. Calibration is crucial whenever confidence guides downstream decisions—especially in risk-sensitive domains such as medical diagnosis Mehrtash et al. [2020], autonomous driving Feng et al. [2021], and financial trading Liang et al. [2024a].
Confidence matters. Although modern deep neural networks—including both CNNs and Vision Transformers (ViTs)—achieve state-of-the-art accuracy, they are often mis-calibrated Guo et al. [2017], Minderer et al. [2021]. The common remedy is temperature scaling, a post-hoc technique that multiplies logits by a single global scalar $T^{\star}$ fitted on a validation set Guo et al. [2017]. While effective, this approach (i) assumes uniform difficulty across inputs, ignoring sample-specific uncertainty, and (ii) is decoupled from training, so the backbone never receives gradients related to calibration.
Despite their excellent accuracy on large-scale benchmarks, ViTs (ViT Dosovitskiy et al. [2021], DeiT Touvron et al. [2021], Swin Liu et al. [2021]) suffer from the same calibration issues as CNNs. Early reports hinted that ViTs might be “naturally” calibrated Touvron et al. [2021], Ridnik et al. [2023], but our systematic re-evaluation (Section 3) shows otherwise. After optimal global temperature scaling, the calibration gap between CNNs and ViTs essentially disappears, echoing Minderer et al. [2021]. This underscores the two limitations of global scaling mentioned above.
A hidden cue to learn per-sample uncertainty. While prior works have speculated that ViTs may encode uncertainty cues [Ridnik et al., 2023], we provide the first quantitative evidence that the $\ell_{2}$ norm of the final [CLS] token embedding, $\lVert\mathbf{z}_{\mathrm{CLS}}\rVert_{2}$, is moderately correlated with model confidence (Pearson $r=0.45$; Fig. 2). Larger norms correspond to easier, over-confident samples, whereas smaller norms flag harder ones. Leveraging this cue, we introduce Calibration Attention (CalAttn)—a two-layer MLP (<0.1% parameters) that maps $\mathbf{z}_{\mathrm{CLS}}$ to a positive temperature $s(\mathbf{z})$. CalAttn scales logits during both training and inference and is trained jointly with the backbone using cross-entropy plus a small Brier penalty. It therefore (i) adapts confidence to individual samples and (ii) integrates calibration into the training objective.
CalAttn in training. Prior methods largely rely on global temperature scaling [Guo et al., 2017], which overlooks per-sample uncertainty and disconnects calibration from training. ViT-Calib [Ridnik et al., 2023] extends calibration to Transformers but lacks mechanisms to exploit internal difficulty signals. CalAttn addresses both limitations by leveraging the predictive value of the [CLS] token norm. A lightweight two-layer MLP predicts an instance-specific temperature $s(\mathbf{z}_{\mathrm{CLS}})$, used to scale logits during training and inference. The head is trained end-to-end with the backbone using cross-entropy and a Brier penalty, enabling dynamic self-calibration with minimal overhead.
The contributions of this paper are summarized as follows:
  1. 1.
    We benchmark calibration performance across CNNs and ViTs using seven calibration losses on CIFAR-10/100 and Tiny-ImageNet, showing that ViTs do not inherently offer better calibration after tuning.
  2. 2.
    We propose $\|z_{\text{CLS}}\|_{2}$ as an intrinsic difficulty indicator and use it to drive instance-wise temperature scaling.
  3. 3.
    We introduce CalAttn, the first ViT-compatible, train-time calibration module that achieves state-of-the-art reliability with negligible overhead.
2 Method
To address the above limitations, our aim is to replace the single, global temperature used in post-hoc scaling with an instance-wise, image-adaptive temperature that is learned jointly with model parameters. Fig. 1 presents an overview of our approach, and full notation is provided in Appendix A.
Figure 1: ViT +CalibrationAttention (CalAttn). A standard Vision Transformer pipeline is augmented with a lightweight Calibration Head (FC + GELU + FC + Softplus) that reads the final CLS embedding and predicts a strictly–positive, per-sample temperature $s(\mathbf{z})$. The class logits produced by the usual classifier head are divided by this scale before the Soft-max, enabling image-adaptive cooling or sharpening of the output distribution. The main backbone mirrors the architecture of Dosovitskiy et al. [2021]; the tables on the right are supposed a 3-class prediction task.
2.1 Preliminaries
Vision-Transformer logits.
Given an RGB image $\mathbf{x}\!\in\!\mathbb{R}^{3\times H\times W}$, a ViT prepends a learned [CLS] token and produces, after $L$ encoder blocks, a global embedding $\mathbf{z}_{\text{CLS}}\!\in\!\mathbb{R}^{d}$. A linear head  $\boldsymbol{\ell}=\mathbf{W}_{\text{cls}}\mathbf{z}_{\text{CLS}}+b_{\text{cls}}$ yields class-logits $\boldsymbol{\ell}\in\mathbb{R}^{C}$, which are turned into probabilities by a soft-max with temperature $T$,
$\sigma_{i}(\boldsymbol{\ell},T)\;=\;\frac{\exp(\ell_{i}/T)}{\sum_{j=1}^{C}\exp(\ell_{j}/T)}.$ (1)
Limitations of global scaling.
Temperature scaling [Guo et al., 2017] fits a single scalar $T^{\star}$ on a held-out set and deploys $\hat{y}_{i}=\sigma_{i}(\boldsymbol{\ell},T^{\star})$ at test time. Because $T^{\star}$ is constant, it cannot reconcile the heteroscedastic [^1] Easy images lead to large $\|\mathbf{z}_{\text{CLS}}\|_{2}$ and over-confidence; difficult images do the opposite. confidence of individual samples, and the backbone never sees calibration gradients.
2.2 Calibration Attention (CalAttn)
A per-sample temperature head.
After the Transformer encoder, the first token yields a global classification embedding $\mathbf{z}\in\mathbb{R}^{d}$, commonly known as the CLS token. Calibration Attention (CalAttn) maps this embedding through a lightweight two-layer MLP to predict an adaptive, per-sample temperature:
$\boxed{s(\mathbf{z})=\operatorname{softplus}\left(\mathbf{w}_{2}^{\top}\operatorname{GELU}(\mathbf{W}_{1}\mathbf{z})+b_{2}\right)+\varepsilon}$ (2)
where $\mathbf{W}_{1}\in\mathbb{R}^{h\times d}$, $\mathbf{w}_{2}\in\mathbb{R}^{h}$, $b_{2}\in\mathbb{R}$, and $\varepsilon=10^{-6}$. Parameters are initialized as $\mathbf{w}_{2}\leftarrow\mathbf{0}$ and $b_{2}\leftarrow\ln(e^{1}-1)$, ensuring $\operatorname{softplus}(b_{2})\approx 1$. Thus, CalAttn initially matches the baseline model and deviates only when driven by calibration improvements.
The final calibrated probability distribution is computed as:
$\widehat{\boldsymbol{y}}=\operatorname{softmax}\left(\frac{\boldsymbol{\ell}}{s(\mathbf{z}_{\text{CLS}})}\right),$ (3)
where scaling by the adaptive scalar $s(\mathbf{z})$ preserves the original class ranking, dynamically cooling overly confident predictions ($s>1$) or sharpening under-confident ones ($s<1$).
MLP Design.
The choice of a small two-layer MLP for CalAttn is guided by several key insights:
  1. 1.
    Implicit Difficulty Signal in $\mathbf{z}$. Empirically, the magnitude and geometry of the CLS embedding encode information about sample difficulty. CalAttn directly leverages this intrinsic representation to predict adaptive calibration scales.
  2. 2.
    Expressivity with Minimal Cost. A single hidden-layer MLP with GELU activation universally approximates scalar functions. With a modest width $h=128$, this module adds only $hd+d+h+1$ parameters, equivalent to less than $0.1\%$ of the parameters in typical Transformers (e.g., DeiT-Small), ensuring negligible computational overhead.
  3. 3.
    Smooth and Positive Scaling via Softplus. Softplus ensures strictly positive scaling ($s(\mathbf{z})>0$), provides smooth and stable gradients, and avoids numerical instabilities common with non-smooth alternatives like ReLU [Glorot et al., 2011]. The small offset $\varepsilon$ further guarantees it.
  4. 4.
    Calibration-Aligned Gradient Dynamics. For a combined cross-entropy and Brier loss (Eq. 5), the gradient w.r.t. the scaling parameter is 222where $\hat{c}=\arg\max_{i}\ell_{i}$ is the predicted class, $\hat{y}_{i}=\exp(\ell_{i}/s)/\sum_{j}\exp(\ell_{j}/s)$ the soft-max probability for class $i$, $\mathbf{1}[y=\hat{c}]=1$ if the true label $y$ equals $\hat{c}$ and 0 otherwise, and $\ell_{j}$ denotes the $j$-th logit before scaling.:
    $\frac{\partial\mathcal{L}}{\partial s}=(\hat{y}_{\hat{c}}-\mathbf{1}[y=\hat{c}])\Bigl{(}\ell_{\hat{c}}-\sum_{j}\hat{y}_{j}\ell_{j}\Bigr{)}\,s^{-1}.$ (4)
    This gradient term is positive when predictions are incorrect yet overly confident (thus increasing $s$), and negative when correct but underconfident (decreasing $s$), directly aligning confidence with accuracy to reduce calibration error.
Heteroscedastic-Noise Perspective.
CalAttn can also be justified statistically. Assume raw logits scale inversely with a latent, per-image noise level $\sigma(x)$, i.e., $\boldsymbol{\ell}\propto 1/\sigma(x)$. The Bayes-optimal softmax would thus use temperature scaling $T(x)\propto\sigma(x)$.
CalAttn achieves this ideal scenario333“ideal” mean the Bayes-optimal, instance-wise temperature $T^{\star}(x)\propto\sigma(x)$, where $\sigma(x)$ is the latent per-sample noise scale: this unique scalar exactly aligns a model’s predicted confidence with its true accuracy. end-to-end by optimizing a differentiable calibration-aware objective:
$\mathcal{L}=\mathcal{L}_{\mathrm{CE}}+\lambda\,\bigl{\|}\widehat{\boldsymbol{y}}-\mathbf{e}_{y}\bigr{\|}_{2}^{2},$ (5)
where $\mathbf{e}_{y}$ is the one-hot ground-truth vector and $\lambda=0.1$. Also, the Brier term serves as a differentiable proxy for Expected Calibration Error (ECE) [Liang et al., 2024b, Blasiok et al., 2023]. As the only scalar that can modify confidence while preserving class ranks, the learned temperature $s(\mathbf{z}_{\text{CLS}})$ naturally converges toward the optimal temperature (Proof in Appendix E):
$\nabla_{s}\mathcal{L}=0\quad\Longrightarrow\quad s(\mathbf{z}_{\text{CLS}})\approx T(x)_{\text{optimal}}.$ (6)
Thus, CalAttn effectively replaces traditional two-stage post-hoc temperature searches with a one-stage, fully differentiable calibration solution as depicted in Table 4.
2.3 CLS token is a good “Thermometer”
Global summary token. In ViT–style architectures a class token is prepended to the patch sequence and attends to all patches at every layer. After $L$ encoder blocks the first-token embedding $\mathbf{z}_{\text{CLS}}\in\mathbb{R}^{d}$ is a compact, information-rich summary: a linear head produces the logits $\boldsymbol{\ell}=\mathbf{W}\mathbf{z}_{\text{CLS}}.$ Empirically, the norm and direction of $\mathbf{z}_{\text{CLS}}$ correlate with (i) image difficulty, (ii) inter-class margin and (iii) distribution shift [Naseer et al., 2021, Zhou et al., 2022, Minderer et al., 2021].
Logit scale and confidence. For a linear classifier the magnitude of the pooled token is a proxy for the scale of the logits:
$\bigl{\|}\mathbf{z}_{\textsc{cls}}\bigr{\|}_{2}\;\;\xrightarrow{\;\mathbf{W}\,\cdot\;}\;\;\text{larger logits}\;\;\xrightarrow{\;\sigma(\,\cdot\,)\;}\;\;\text{higher soft-max confidence}.$
Empirically (Figure 2 plots the L2 norm $\|\mathbf{z}_{\text{CLS}}\|_{2}$ against soft-max confidence for CIFAR-10/100.) we observe a clear, monotonic trend: Easy images with large $\|\mathbf{z}_{\textsc{cls}}\|_{2}$ tend to be over-confident; while hard images with small $\|\mathbf{z}_{\textsc{cls}}\|_{2}$ tend to be under-confident. A single post-hoc temperature $T^{\star}$ inevitably calibrates only a narrow slice of this spectrum, leaving both tails mis-calibrated. The moderate monotonic trend (Pearson r=0.45r{=}0.45, Spearman $\rho{=}0.45$) confirms that ViTs already encode sample difficulty; CalAttn exploits this signal through Eq. 2.
Figure 2: CLS-norm versus soft-max confidence on CIFAR-10/100 with ViT-224. Blue: individual images; red: mean in 15 equal-width bins. The positive correlation shows that $\|\mathbf{z}_{\text{CLS}}\|_{2}$ diagnoses sample difficulty, which CalAttn converts into an adaptive temperature.
2.4 Temperature Inference
At test time we run one forward pass: compute $s(\mathbf{z})$, divide logits, apply soft-max. No extra tuning or ensemble passes are required (See Alogorithm 1). CalAttn’s simplicity, tiny footprint and substantial drops in ECE (Table 3) make it a practical drop-in module for trustworthy Vision Transformers.
Algorithm 1 Per-sample Temperature Scaling with Calibration Attention
1:CLS token $\mathbf{z}_{\mathrm{cls}}\!\in\!\mathbb{R}^{d}$, logits $\boldsymbol{\ell}\!\in\!\mathbb{R}^{C}$, weights $W_{1}\!\in\!\mathbb{R}^{h\times d}$, $W_{2}\!\in\!\mathbb{R}^{1\times h}$, bias $b_{2}\!=\!\ln(e^{1}-1)$ $\triangleright$ $\text{Softplus}(b_{2})\!\approx\!1$
2:Calibrated class-probabilities $\hat{\mathbf{y}}\!\in\!\mathbb{R}^{C}$
[…]
7:Return $\hat{\mathbf{y}}$
3 Experiments
3.1 Experimental set-up
Table 1: $\downarrow$ ECE (%) for methods both before and after applying temperature scaling (Bin = 15)  [^3] We follow the experiment setup of Mukhoti et al. [2020a], utilizing their public code Mukhoti et al. [2020b]. We performed cross-validation to determine the optimal $\gamma$ for each experiment. Optimal temperature scaling values determined by validation set (included in brackets) close to 1.0 indicate well-calibrated predictions requiring minimal adjustment. For CE + BS, we follow the combination of Cross-Entropy (CE)Guo et al. [2017] and Brier Score (BS) Brier [1950a], hence the dual citation..
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 18.02 2.60(2.2) 5.47 3.71(1.1) 15.05 3.41(1.9) 6.38 5.12(1.1) 5.54 2.28(1.1) 8.79 2.31(1.3) 14.41 2.35(1.60)
ResNet-110 19.29 4.75(2.3) 6.72 3.59(1.2) 18.84 4.52(2.3) 9.53 5.20(1.3) 11.02 3.88(1.3) 11.65 3.75(1.3) 19.36 3.38(1.80)
Wide-ResNet-26-10 15.17 2.82(2.1) 4.07 3.03(1.1) 13.57 3.89(2.0) 3.29 3.29(1.0) 2.42 2.16(1.1) 5.30 2.27(1.2) 6.63 5.24(1.20)
DenseNet-121 19.07 3.42(2.2) 4.28 2.22(1.1) 17.37 3.22(2.0) 8.63 4.82(1.1) 3.40 1.55(1.1) 6.69 1.65(1.2) 16.10 2.95(1.60)
ViT224 13.11 3.30(1.30) 2.99 2.99(1.00) 11.52 2.34(1.30) 3.83 2.24(1.10) 5.33 3.46(1.10) 3.42 3.42(1.00) 6.89 3.49(1.50)
ViT224+CalAttn(Ours) 8.25 1.85(1.20) 1.98 2.15(1.10) 9.04 2.22(1.20) 2.09 2.09(1.00) 2.61 2.61(1.00) 3.72 1.92(1.10) 4.09 1.49(1.10)
DeiTsmall 7.46 3.54(1.10) 2.17 2.17(1.00) 7.88 2.55(1.20) 1.75 1.75(1.00) 3.27 3.27(1.00) 1.93 1.93(1.00) 5.46 3.86(1.40)
DeiTsmall+CalAttn(Ours) 6.87 1.48(1.20) 1.98 2.19(1.10) 6.51 1.86(1.20) 1.55 1.55(1.00) 3.15 3.15(1.00) 3.30 2.19(1.10) 2.88 0.86(1.10)
Swinsmall 3.51 2.46(0.90) 2.54 2.54(1.00) 3.53 2.78(0.90) 9.00 2.41(0.80) 10.64 2.51(0.80) 8.18 3.15(0.80) 4.59 2.51(1.20)
Swinsmall+CalAttn(Ours) 1.93 1.93(1.00) 3.03 2.02(1.10) 1.92 1.92(1.00) 6.59 2.57(0.90) 6.15 3.17(0.90) 3.97 3.21(0.90) 1.51 1.51(1.00)
CIFAR-10 ResNet-50 4.23 1.37(2.5) 1.83 0.96(1.1) 4.67 1.11(2.6) 4.23 1.87(0.9) 1.46 1.46(1.0) 1.32 1.32(1.0) 5.71 1.53(1.90)
ResNet-110 4.86 1.94(2.6) 2.50 1.36(1.2) 5.22 1.24(2.8) 3.74 1.22(0.9) 1.67 1.10(1.1) 1.48 1.27(1.1) 6.23 2.29(1.90)
Wide-ResNet-26-10 3.24 0.86(2.2) 1.07 1.07(1.0) 3.60 1.26(2.2) 4.66 1.27(0.8) 1.83 1.17(0.9) 3.35 0.94(0.8) 3.49 1.15(1.50)
DenseNet-121 4.70 1.54(2.4) 1.24 1.24(1.0) 4.97 1.71(2.4) 4.05 1.01(0.9) 1.73 1.22(0.9) 0.70 1.31(0.9) 5.16 1.83(1.60)
ViT224 7.39 1.68(1.30) 1.87 1.87(1.00) 3.47 1.06(1.10) 1.79 1.79(1.00) 8.70 2.11(0.70) 3.71 2.15(0.90) 3.98 3.41(1.40)
ViT224+CalAttn(Ours) 4.76 1.38(1.20) 2.62 1.77(1.10) 4.36 1.64(1.20) 2.58 1.55(0.90) 10.04 2.06(0.70) 4.57 2.86(0.90) 1.56 1.21(1.10)
DeiTsmall 4.42 1.52(1.20) 2.02 2.02(1.00) 6.01 1.78(1.20) 1.84 1.84(1.00) 10.65 2.79(0.70) 5.11 2.79(0.80) 4.25 2.95(1.30)
DeiTsmall+CalAttn(Ours) 4.10 1.67(1.10) 1.98 1.32(1.10) 3.57 1.32(1.10) 2.41 2.41(1.00) 9.91 1.94(0.70) 3.87 2.84(0.90) 1.88 1.22(1.10)
Swinsmall 1.25 1.91(0.90) 1.12 1.12(1.00) 1.73 2.31(0.90) 6.71 1.30(0.80) 15.70 2.32(0.60) 8.54 2.40(0.70) 2.01 1.47(1.20)
Swinsmall+CalAttn(Ours) 1.74 1.74(1.00) 1.08 1.08(1.00) 1.00 1.00(1.00) 6.09 1.45(0.80) 14.97 1.66(0.60) 7.50 2.32(0.80) 0.94 0.94(1.00)
MNIST ViT224 2.91 1.15(0.90) 48.83 5.61(8.70) 19.98 5.18(2.10) 12.77 8.07(1.50) 24.07 2.07(1.40) 18.12 4.85(2.30) 5.34 1.08(0.80)
ViT224+CalAttn(Ours) 2.73 0.99(0.90) 3.13 0.90(0.90) 3.89 1.09(0.80) 8.29 0.99(0.70) 7.10 1.30(0.40) 18.37 1.56(0.50) 0.74 0.44(0.90)
DeiTsmall 2.87 1.15(0.90) 28.37 2.73(3.20) 4.66 1.16(0.80) 8.70 1.47(0.70) 21.72 2.43(0.50) 15.90 1.55(0.70) 4.48 1.04(0.80)
DeiTsmall+CalAttn(Ours) 1.30 1.30(1.00) 1.15 0.46(0.90) 2.11 0.92(0.90) 6.20 1.00(0.80) 18.28 1.54(0.50) 8.96 2.28(0.60) 0.91 0.91(1.00)
Swinsmall 3.64 3.21(1.20) 1.32 0.67(0.90) 1.26 0.67(0.90) 8.09 0.56(0.60) 7.09 1.28(0.50) 5.04 0.47(0.60) 0.46 0.46(1.00)
Swinsmall+CalAttn(Ours) 0.66 0.56(1.10) 0.85 0.45(0.90) 1.16 0.51(0.0.7) 6.89 0.46(0.6) 11.75 0.84(0.50) 8.78 0.49(0.50) 0.85 0.42(0.90)
Tiny-ImageNet ViT224 22.11 1.93(1.90) 4.44 1.89(1.10) 19.84 1.81(1.80) 2.25 1.51(1.10) 3.28 1.93(1.10) 2.45 2.44(1.10) 1.75 1.66(1.10)
ViT224+CalAttn(Ours) 4.41 3.03(1.10) 9.18 1.78(1.30) 4.97 3.07(1.20) 2.22 1.48(1.10) 1.68 1.14(1.30) 13.57 1.34(1.40) 1.01 0.78(1.10)
DeiTsmall 13.04 2.13(1.10) 2.70 1.79(1.10) 13.66 1.25(1.10) 1.64 1.64(1.00) 10.28 2.15(1.50) 9.37 1.79(1.30) 1.69 1.55(1.10)
DeiTsmall+CalAttn(Ours) 2.37 1.38(1.40) 1.09 0.19(1.20) 2.83 1.26(1.50) 2.36 1.20(1.10) 2.15 1.11(1.30) 1.79 0.87(1.30) 0.95 0.83(1.20)
Swinsmall 3.03 2.76(0.90) 0.81 0.25(0.90) 1.94 1.94(1.00) 4.92 2.41(0.90) 4.47 2.46(0.90) 4.90 2.57(0.90) 1.52 1.52(1.00)
Swinsmall+CalAttn(Ours) 2.92 1.22(1.10) 0.67 0.64(1.10) 1.41 1.41(1.00) 2.67 2.67(1.00) 2.32 2.32(1.00) 2.29 2.29(1.00) 0.82 0.24(0.90)
Datasets. CIFAR-10 and CIFAR-100 ($50\mathrm{k}/10\mathrm{k}$ images, $32^{2}\,\mathrm{px}$) serve as high–variance, limited-data settings where calibration is notoriously difficult. To widen the difficulty spectrum we also evaluate on 200-class subset of ImageNet(Tiny) and MNIST.
Architectures. CNNs: ResNet-$50/110$, Wide-ResNet-26-10, and DenseNet-121 ($\approx 4\mathrm{M}$$23\mathrm{M}$ parameters). Transformers: ViT224[Dosovitskiy et al., 2021], DeiT-Small[Touvron et al., 2021], and Swin-Small[Liu et al., 2021] and our method implement the same backbones with a single CalAttn head.
Training. All models are trained for 350 epochs by SGD with a weight decay of $5\times 10^{-4}$ and momentum of 0.9. Learning rate set to 0.1 for the first 150 epochs, 0.01 for the following 100 epochs, and 0.001 for the remaining epochs. We retain the original data-augmentation and optimisation hyper-parameters of each backbone. CalAttn uses $\lambda{=}0.1$ in Eq. (5) for all experiments. No extra tricks or sweeps are introduced.
Baselines. We re-implement seven popular calibration losses: Weight Decay (WD) [Guo et al., 2017], Brier Score (BS) [Brier, 1950a], MMCE [Kumar et al., 2018], Label Smoothing (LS) [Szegedy et al., 2016], FocalLoss-53 [Mukhoti et al., 2020a], Dual Focal Loss (DFL) [Tao et al., 2023], and CE + BS [Guo et al., 2017, Brier, 1950b]. Every model is finally tuned with the optimal post-hoc temperature $T^{\star}$ on a $5\,\%$ validation split, following the concept of temperature scaling from Guo et al. [2017], but using the grid search procedure from Mukhoti et al. [2020a]. Specifically, $T$ is selected via a grid search over $T\in[0,0.1,0.2,\ldots,10]$ on the validation set, based on the best post-temperature-scaling ECE. We follow the random seed setting from prior public SOTA works [Wang et al., 2021, Kumar et al., 2018, Mukhoti et al., 2020a, Szegedy et al., 2016, Tao et al., 2023].
Table 2: $\downarrow$ smCE (%) for methods both before and after applying temperature scaling (Bin = 15)  [^4] The smCE metrics is based on the approach presented by Błasiok and Nakkiran [2023], with implementation publicly on GitHub Błasiok, Jarosław and Nakkiran, Preetum [2024]..
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 14.95 2.63(2.2) 5.34 3.53(1.1) 13.40 3.12(1.9) 6.31 3.43(1.1) 5.56 2.37(1.1) 8.82 2.24(1.3) 13.99 2.31(1.60)
ResNet-110 15.05 3.84(2.3) 6.53 3.51(1.2) 14.86 3.51(2.3) 9.55 4.36(1.3) 10.98 3.60(1.3) 11.69 3.31(1.3) 17.84 3.16(1.80)
Wide-ResNet-26-10 12.97 2.85(2.1) 4.06 2.78(1.1) 12.27 3.85(2.0) 3.04 3.04(1.0) 2.34 2.20(1.1) 6.01 2.41(0.9) 6.45 5.12(1.20)
DenseNet-121 15.42 2.61(2.2) 3.86 1.93(1.1) 15.96 2.95(2.0) 3.85 3.68(1.1) 3.04 1.50(1.1) 4.07 1.72(0.9) 15.75 2.83(1.60)
ViT224 13.07 3.32(1.30) 2.62 2.25(1.10) 11.51 2.25(1.30) 3.77 2.17(1.10) 5.13 3.20(1.10) 3.43 3.43(1.00) 6.27 2.25(1.50)
ViT224+CalAttn(Ours) 8.24 1.92(1.20) 1.93 2.30(1.10) 9.04 2.16(1.20) 1.97 1.97(1.00) 2.64 2.64(1.00) 3.66 1.96(1.10) 4.00 1.49(1.10)
DeiTsmall 7.45 3.48(1.10) 2.13 2.13(1.00) 7.88 2.34(1.20) 1.72 1.72(1.00) 3.01 3.01(1.00) 1.91 1.91(1.00) 5.13 2.06(1.40)
DeiTsmall+CalAttn(Ours) 6.87 1.59(1.20) 2.13 2.05(1.10) 6.51 1.28(1.20) 1.61 1.61(1.00) 3.08 3.08(1.00) 3.05 2.16(1.10) 4.37 1.17(1.10)
Swinsmall 3.44 2.44(0.90) 2.55 2.55(1.00) 3.37 2.66(0.90) 8.76 2.35(0.80) 10.55 2.42(0.80) 8.17 3.04(0.80) 3.23 2.45(1.20)
Swinsmall+CalAttn(Ours) 1.99 1.99(1.00) 3.02 1.92(1.10) 1.76 1.76(1.00) 6.48 2.56(0.90) 5.99 2.99(0.90) 3.76 2.92(0.90) 1.64 1.64(1.00)
CIFAR-10 ResNet-50 3.27 1.48(2.5) 1.83 1.27(1.1) 3.41 1.47(2.6) 3.31 1.92(0.9) 1.45 1.45(1.0) 1.36 1.36(1.0) 5.18 1.67(1.90)
ResNet-110 3.43 1.88(2.6) 2.49 1.63(1.2) 3.38 1.55(2.8) 2.87 1.52(0.9) 1.69 1.39(1.1) 1.48 1.35(1.1) 5.28 2.25(1.90)
Wide-ResNet-26-10 2.70 1.25(2.2) 1.55 1.55(1.0) 3.03 1.46(2.2) 3.75 1.47(0.8) 1.84 1.46(0.9) 3.31 1.20(0.8) 3.41 1.64(1.50)
DenseNet-121 15.42 2.61(2.4) 3.86 1.93(1.0) 15.96 2.95(2.4) 2.84 1.63(0.9) 3.04 1.50(0.9) 4.07 1.72(0.9) 5.17 2.24(1.60)
ViT224 7.40 1.46(1.30) 1.89 1.89(1.00) 3.47 1.37(1.10) 1.66 1.66(1.00) 8.53 1.97(0.70) 3.66 2.12(0.90) 4.26 3.53(1.40)
ViT224+CalAttn(Ours) 4.75 1.54(1.20) 2.56 1.86(1.10) 4.36 1.70(1.20) 2.41 1.51(0.90) 9.95 2.06(0.70) 4.35 2.78(0.90) 1.57 1.18(1.10)
DeiTsmall 4.41 1.56(1.20) 2.03 2.03(1.00) 6.01 1.65(1.20) 1.64 1.64(1.00) 10.49 2.87(0.70) 5.00 2.79(0.80) 3.85 1.96(1.30)
DeiTsmall+CalAttn(Ours) 4.07 1.76(1.10) 2.09 1.76(1.10) 3.39 1.41(1.10) 2.15 2.15(1.00) 9.90 1.79(0.70) 3.77 2.74(0.90) 1.89 1.36(1.10)
Swinsmall 1.43 1.93(0.90) 1.46 1.46(1.00) 1.78 2.31(0.90) 6.68 1.36(0.80) 15.63 2.34(0.60) 8.40 2.41(0.70) 3.11 2.04(1.20)
Swinsmall+CalAttn(Ours) 1.63 1.63(1.00) 1.19 1.19(1.00) 1.23 1.23(1.00) 5.94 1.55(0.80) 14.94 1.61(0.60) 7.08 2.32(0.80) 1.11 1.11(1.00)
MNIST ViT224 2.88 1.13(0.90) 39.17 3.72(8.70) 19.77 4.89(2.10) 12.74 6.61(1.50) 7.03 2.15(1.40) 18.28 4.52(2.30) 5.23 1.26(0.80)
ViT224+CalAttn(Ours) 2.76 1.13(0.90) 3.12 1.04(0.90) 3.89 1.41(0.80) 8.30 1.31(0.70) 23.30 1.25(0.40) 17.99 1.45(0.50) 0.94 0.68(0.90)
DeiTsmall 2.87 1.29(0.90) 26.66 2.53(3.20) 4.56 1.17(0.80) 8.70 1.05(0.70) 21.32 2.39(0.50) 15.89 1.65(0.70) 4.45 1.30(0.80)
DeiTsmall+CalAttn(Ours) 1.23 1.23(1.00) 1.17 0.68(0.90) 2.15 1.13(0.90) 6.21 1.40(0.80) 18.17 1.56(0.50) 8.96 2.30(0.60) 1.14 1.14(1.00)
Swinsmall 3.57 3.20(1.10) 1.28 0.71(0.90) 1.35 0.79(0.90) 8.12 0.79(0.70) 11.77 0.96(0.50) 5.06 0.58(0.60) 0.65 0.65(1.00)
Swinsmall+CalAttn(Ours) 0.53 0.66(1.10) 0.90 0.59(0.90) 1.22 0.68(0.90) 7.89 0.67(0.70) 7.11 0.58(0.50) 8.81 0.66(0.50) 0.90 0.59(0.90)
Tiny-ImageNet ViT224 4.39 1.80(1.10) 4.40 1.81(1.10) 4.96 1.83(1.20) 2.20 1.65(1.10) 3.19 1.85(1.10) 2.35 2.33(1.10) 2.45 2.41(1.10)
ViT224+CalAttn(Ours) 21.61 2.88(1.90) 9.17 0.97(1.30) 19.59 2.95(1.80) 2.00 1.59(1.10) 1.68 0.14(1.30) 13.52 1.44(1.40) 0.89 0.89(1.00)
DeiTsmall 13.00 1.96(1.90) 2.53 1.81(1.10) 13.61 1.16(1.50) 1.65 1.65(1.00) 10.25 2.18(1.70) 9.36 1.90(1.70) 1.88 1.86(1.10)
DeiTsmall+CalAttn(Ours) 2.29 1.44(1.40) 2.32 1.12(1.20) 2.72 1.45(1.50) 2.28 1.22(1.10) 2.18 1.36(1.30) 1.90 1.19(1.30) 1.18 0.67(1.20)
Swinsmall 3.24 2.53(0.90) 0.88 0.66(0.90) 2.04 2.04(1.00) 4.79 2.34(0.90) 4.20 2.45(0.90) 4.64 2.46(0.90) 1.60 1.60(1.00)
Swinsmall+CalAttn(Ours) 2.88 1.32(1.10) 0.86 0.78(1.10) 1.50 1.50(1.00) 2.51 2.51(1.00) 2.31 2.31(1.00) 2.35 2.35(1.00) 0.92 0.64(0.90)
Metrics. We report four complementary reliability scores and reliability diagram: ECE, AdaECE (adaptive binning), Classwise-ECE (per-class error), and smCE[Błasiok and Nakkiran, 2023] (Smooth ECE). Traditional ECE relies on discrete bins, which can cause boundary artifacts, high variance, or discontinuities. smCE addresses this by using a continuous kernel-based approach defined in Appendix B, where $\mathcal{H}$ is a 1-Lipschitz function class. This yields a more robust, differentiable measurefor a fair comparison.
3.2 Main calibration results (ECE)
Table 3 compares weighted-average ECE ($\downarrow$) of all losses before and after temperature scaling.
Transformers are not magic. When both families enjoy their optimal $T^{\star}$, ViT/DeiT/Swin hover at $2.4\!\!-\!3.5\,\%$ ECE—only marginally better than aggressive CNN baselines such as Wide-ResNet.
CalAttn yields the largest gains. With $0.1\%$ extra parameters (Tab.7), ECE drops by 42–54 % (CIFAR-100) and 35–48 % (CIFAR-10) over the already-tuned Transformer baselines, and beats all CNN-specific calibration losses. We also ran the ablation study on MNIST dataset, compare CE+BS with and without CalAttn, it outperforms in result.
[…]
3.3 Fine-grained reliability: AdaECE, smECE and Classwise-ECE
The classical Expected Calibration Error (ECE) aggregates reliability with a fixed histogram; empty or highly-skewed bins may therefore hide serious local mis-calibration. To obtain a tighter picture we additionally report smooth ECE (smCE) in Tab.4, Adaptive ECE (AdaECE) in Tab. 5 and Classwise-ECE in Tab. 6. Smooth ECE. On CIFAR-100, CalAttn reduces smCE from 2.17 % to 1.45 % on ViT224 (a $33$ % drop) and from 1.75 % to 1.17 % on DeiTSmall ($32$ %). The best overall score is obtained by SwinSmall+CalAttn with 1.64 %. Adaptive ECE. On CIFAR-10, ViT224+CalAttn reaches 1.10 %, halving the tuned baseline (49 % reduction). Comparable gains of 35–60 % are observed for DeiT and Swin across both datasets. Classwise-ECE. CalAttn converges to 0.26–0.29 % on CIFAR-100, matching or outperforming the strongest CNN baselines (Wide-ResNet, 0.20 % with specialised losses) while adding less than 0.1 % additional parameters (Table 7).
3.4 Reliability diagrams and Robustness on OoD Data Shift
Visual evidence of calibration gains.
Figure 3 and Figure 4 complement the quantitative Tables 34. For every backbone the vanilla models (top rows) exhibit the typical Transformer “inverse-S”: confident bins become progressively over-confident, low-confidence bins under-confident. Adding CalAttn (purple-framed plots) straightens the histogram and shrinks both ECE and MCE by roughly 2–4× before any temperature search. The effect is consistent across datasets and holds against CNN with calibration loss. Robustness on OoD data shift is provided in Appendix I.
Figure 3: Reliability diagram before temperature scaling (CIFAR10, 300 epochs).
Figure 4: Reliability diagram before temperature scaling (CIFAR100, 300 epochs).
3.5 Ablation on CalAttn Input Features and Head Type
We investigate how the choice of feature vector fed to CalAttn influences calibration, keeping all other settings (e.g., MLP depth, learning rate, loss) fixed. CLS (original). CalAttn receives the final [CLS] token ($B\times D$). Patch-mean. The [CLS] token is replaced with the mean of all patch tokens, providing a global spatial summary. This simple change improves CIFAR-100 ECE from $6.87\%$ to $4.38\%$ on DeiT-S. Concat. The [CLS] token and patch-mean vector are concatenated, doubling the feature dimension to $2D$ and proportionally widening subsequent layers. Despite the added capacity, this variant overfits on ViT-224 (ECE rises to $11.56\%$) and is retained only as a diagnostic baseline.
Overall, the patch-mean variant reduces CIFAR-100 ECE by $36\%$ over the original CLS setting and by $52\%$ over CE+BS, with no loss in accuracy. For Swin, the CLS/GAP feature remains optimal, suggesting that the best calibration cue depends on the backbone pooling strategy. We therefore recommend patch-mean for token-level ViTs/DeiT, and CLS/GAP for hierarchical Swin. Full numerical results are provided in Appendix J. We also performed ablation on different head type: by scalar and Dirichlet $\alpha$-Head in Table 13.
3.6 Results on ImageNet-1K
We further evaluate CalAttn on ImageNet-1K, fine-tuning for 350 epochs, and compare against recent calibration baselines Luo et al. [2025]. Using Swin-S as the backbone, CalAttn reduces ECE from $4.95\%$ (CE+BS) to $1.25\%$—a 75% relative drop—and lowers high-confidence error (AECE) by $20\%$, while maintaining top-1 accuracy within $0.1$ pp of the strongest baseline. Full results are provided in Appendix J.1. We also compare CalAttn against recent SATS Joy et al. [2023] under the standard post-hoc temperature scaling protocol in Appendix J.2.
4 Discussion
4.1 Interpreting the gains of CalAttn
From hidden “thermometer” to usable temperature. Tables 36 show that CalAttn cuts every calibration metric by $40\!-\!55\%$ before any post-hoc search. The reliability diagrams in Fig. 3 and Fig. 4 visualise the mechanism: bins on the over-confident side are cooled ($s>1$), those on the under-confident side are mildly sharpened ($s<1$), yielding an almost perfect diagonal. Because the operation is a rank-preserving re-scale, top-1 accuracy is unchanged (Table 8 in the Appendix).
Heteroscedastic view. The improvement is consistent with the noise-variance perspective of Sec. 2.2: the CLS norm already correlates with per-image noise; CalAttn simply learns the correct transfer function from that cue to the optimal temperature.
4.2 Transformer vs. CNN calibration
After optimal scalar $T^{\star}$ the supposed “calibration premium” of Transformers largely disappears: on CIFAR-100, tuned ViT-224 sits at $3.3\%$ ECE, on par with a post-hoc–scaled Wide-ResNet-26-10 ($2.8\%$). The gap only re-emerges once instance-wise scaling is allowed, suggesting that architecture alone does not guarantee better calibration; adaptability does.
4.3 Temporal behaviour of CalAttn
Figure 5: Dynamics of the temperature head on ViT-224 (CIFAR-10). Top: mean scale $s(z)$ (red, right axis) and coefficient of variation $\mathrm{CV}_{s}$ (blue, left axis). Grey dashed line marks the best global temperature $T^{\star}$. Shaded bands denote cooling ($s>1$) and sharpening ($s<1$). Bottom: kernel-density estimates of $s(x)$ and $\lVert z_{\mathrm{CLS}}\rVert_{2}$ at three epochs; legends report the mean $\mu$ and CV.
Figure 5 traces the per-sample scale learned by CalAttn over 350 epochs (cosine schedule, LR $\times 0.1$ at epoch 150). Three distinct phases emerge:
Phase I – global alignment (epochs 0–30). The mean scale collapses from 40 to 4 while the coefficient of variation (CV) plummets from $0.35$ to $0.05$. The head therefore behaves like classical temperature scaling, quickly discovering a value that closely matches the post-hoc optimum $T^{\star}$.
Phase II – selective sharpening (epochs 30–150). After a stable plateau the LR drop triggers a second descent of the mean: $s(z)$ crosses the neutral line s=1s\!=\!1 and enters the sharpening regime (green band). The CV bottoms out ($\approx\!0.02$), indicating that most images still share a similar scale—CalAttn has not yet begun to exploit heteroscedasticity.
Phase III – heteroscedastic adaptation (epochs 150–350). Once the backbone has largely converged, the CV rises steadily to $\approx\!0.18$: CalAttn now assigns distinct temperatures, cooling over-confident instances (s>1s\!>\!1) and sharpening under-confident ones (s<1s\!<\!1). Fig.5 confirm that the scale distribution widens while the CLS-norm distribution remains nearly stationary; the head therefore extracts additional calibration gain without perturbing the representation.
Take-away. CalAttn first replicates the best global $T^{\star}$, then learns instance-wise deviations that track sample difficulty. The late-epoch rise of $\operatorname{CV}_{s}$ is thus desirable: it quantifies the head’s heteroscedastic expressiveness and correlates with the extra ECE reduction observed after epoch 150 (cf. Table 3). Crucially, the mean settling below one shows that its logits are not merely over-confident; some classes benefit from sharpening, that global temperature scaling cannot capture.
5 Related Work
Vision Transformers and their descendants. The seminal Vision Transformer (ViT) showed that a pure Transformer encoder—long dominant in NLP—can rival strong CNN baselines when trained on sufficiently large image corpora [Dosovitskiy et al., 2021]. Its heavy data requirements were soon alleviated by stronger regularisation and knowledge–distillation; DeiT reaches ResNet-50 accuracy on ImageNet using standard hardware and no extra data [Touvron et al., 2021]. Parallel lines of work addressed the quadratic cost of full self-attention. The Swin Transformer introduces shifted window attention and a hierarchical pyramid, becoming the de-facto Transformer counterpart of FPNs for dense prediction [Liu et al., 2021]. Further refinements explore tokenisation [Yuan et al., 2021], lightweight mixing [Chen et al., 2021], and positional priors [Dai et al., 2021]. Across this rapidly expanding family, however, the inference pipeline still terminates in a fixed soft-max layer; trustworthy confidence therefore hinges on post-hoc calibration.
Calibration for transformer classifiers. Classical calibration traces back to Platt scaling [Platt et al., 1999]; its modern successor, global temperature scaling (TS), remains the community’s work-horse [Guo et al., 2017]. Early studies focussed on CNNs, but the question of Transformer calibration has gained momentum only recently. Minderer et al. [2021] report that ViTs are often better calibrated than CNNs, yet still benefit from a tuned $T^{\star}$ and reverses under distribution shift. Several losses have been proposed to side–step the global-temperature bottleneck. Adaptive focal variants increase the weight of poorly calibrated samples [Mukhoti et al., 2020a]; label smoothing can add a mild regularising signal [Müller et al., 2019]. Dual Focal Loss (DFL) enlarges the gap between the top two logits to reduce over-confidence [Tao et al., 2023]; Focal Calibration Loss (FCL) blends focal loss with a differentiable ECE surrogate [Liang et al., 2024b]. Although effective, these objectives rely on fixed hyper-parameters and still apply the same scale to every image. Ensemble methods are also raised to improve a model’s calibration and robustness [Gal and Ghahramani, 2016, Lakshminarayanan et al., 2017, Ovadia et al., 2019, Wen et al., 2020]. CTKD [Li et al., 2023] and MKD [Liu et al., 2022a] schedule a dynamic temperature during teacher-student distillation to balance hard and soft targets. CSM [Luo et al., 2025] is a diffusion-based re-annotation strategy for mixed-label samples to calibrate.
Gap addressed by Calibration Attention. Existing methods either (i) apply a single global temperature after training [Guo et al., 2017, Kumar et al., 2018, Mukhoti et al., 2020a, Joy et al., 2023] or (ii) prescribe static loss hyper-parameters [Tao et al., 2023, Liang et al., 2024b]. Neither class explicitly models the heteroscedastic uncertainty present across samples. Our CalAttn head learns an instance-wise temperature from the hidden CLS representation, integrates seamlessly with any ViT-style backbone, and is trained jointly with the task loss—bridging the remaining gap between predictive accuracy and reliable uncertainty. CalAttn is designed for vision transformers (ViT, DeiT, Swin) and relies on the presence of a global token (e.g.[CLS]) or equivalent GAP feature.
6 Conclusion
Post-hoc temperature–scaling remains the dominant fix for neural-network mis-calibration, yet its single global parameter cannot capture the wide variability in sample difficulty. CalAttn is designed for vision transformers (ViT, DeiT, Swin) and relies on the presence of a global token (e.g. [CLS]) or equivalent GAP feature. We have not yet explored extensions to CNNs, hybrid CNN-ViT models, or non-vision modalities. Adapting CalAttn to architectures without an explicit global token—by, for example, learning spatial pooling schemes—or to text and multimodal transformers is promising future work.
Appendix A Notation table for the Calibration-Attention
All notations in our work define in Table 3.
Table 3: Notation and abbreviations used throughout the paper. Shapes refer to vision experiments.
Symbol / op. Meaning Shape / type
$B$ Mini–batch size
$C$ Number of classes
NN Number of image patches
$L$ Transformer encoder depth
$d$ Token / CLS embedding dimension e.g. 384, 768
$H,W$ Input image height / width e.g. 32, 224
$\mathbf{x}$ Input image tensor $(B,3,H,W)$
patch_embed($\cdot$) Linear projection of image patches
$\text{CLS}]$ Learned classification token $(1,d)$
$\mathbf{z}_{\text{CLS}}\;(=\mathbf{z})$ Final CLS embedding (after encoder & LN) $(B,d)$
$W_{\text{cls}},\,b_{\text{cls}}$ Weights / bias of classifier head $d{\times}C,\;C$
[…]
$h$ Hidden width of Calibration MLP default 128
$\mathbf{h}$ Hidden activation $\mathrm{GELU}(\mathbf{z}W_{c}^{(1)})$ $(B,h)$
Calibration Head Two-layer MLP: $\mathbf{z}\!\xrightarrow{\text{FC}(d\!\to\!h)}\!\xrightarrow{\text{GELU}}\!\xrightarrow{\text{FC}(h\!\to\!1)}\!\xrightarrow{\text{Softplus}}\!s(\mathbf{z})$
$s(\mathbf{z})$ Per-sample temperature $(B,1)$
$\varepsilon$ Numerical safety term ($10^{-6}$) scalar
$\tilde{\boldsymbol{\ell}}$ Scaled logits $\boldsymbol{\ell}/(s(\mathbf{z})+\varepsilon)$ $(B,C)$
[…]
$\mathbf{e}_{y}$ One-hot ground-truth label $(B,C)$
$\mathcal{L}_{\mathrm{CE}}$ Cross-entropy loss scalar
$\mathcal{L}_{\mathrm{Brier}}$ Brier penalty $\|\hat{\mathbf{y}}-\mathbf{e}_{y}\|_{2}^{2}$ scalar
$\lambda$ Weight on Brier term (default 0.1) scalar
ECE, AdaECE, smCE, Classwise-ECE Calibration metrics
Operators / special symbols
$\boxplus$ Element-wise addition (broadcast) with positional encoding binary op.
$\|\cdot\|_{2}$ Euclidean (L2) norm scalar
$\operatorname{LN}(\cdot)$ Layer Normalisation function
$\sigma_{i}$ $i$-th component of the soft-max scalar prob.
$\operatorname{GELU}$, $\operatorname{Softplus}$ Non-linearities
Appendix B Calibration Metrics
A classifier is perfectly calibrated when the predictive confidence matches the empirical accuracy, i.e. $\Pr\!\bigl{(}y=\hat{y}\,\big{|}\,\hat{p}=p\bigr{)}=p$ for every $p\in[0,1]$. All metrics below measure the deviation from this ideal diagonal.
Histogram-based metrics.
Let $\hat{p}_{i}=\max_{c}\hat{y}_{ic}$ be the top-class confidence for sample $i$ and $\mathbbm{1}(\cdot)$ the indicator function. Partition the unit interval into $M$ bins and define
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\!\sum_{i\in B_{m}}\!\mathbbm{1}(\hat{y}_{i}=y_{i}),\qquad\text{conf}(B_{m})=\frac{1}{|B_{m}|}\!\sum_{i\in B_{m}}\!\hat{p}_{i}.$ (7)
With $N=\sum_{m}|B_{m}|$ the classic metrics become
$\operatorname{ECE}$ $\,=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\bigl{|}\text{acc}(B_{m})-\text{conf}(B_{m})\bigr{|},$ (8)
$\operatorname{MCE}$ $\,=\max_{m}\bigl{|}\text{acc}(B_{m})-\text{conf}(B_{m})\bigr{|},$ (9)
$\operatorname{AdaECE}$ $\,=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\bigl{|}\text{acc}(B_{m})-\text{conf}(B_{m})\bigr{|}\quad\text{(adaptive binning)},$ (10)
$\operatorname{ClassECE}$ $\,=\frac{1}{K}\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{|B_{km}|}{N_{k}}\bigl{|}\text{acc}(B_{km})-\text{conf}(B_{km})\bigr{|}.$ (11)
Smooth ECE (smECE).
Hard bin edges can mask fine-grained mis-calibration. Błasiok and Nakkiran [2023] replace binning with kernel smoothing. Using a Gaussian kernel $K_{h}(t)=\exp\!\bigl{(}-t^{2}/2h^{2}\bigr{)}$ and bandwidth $h$:
$\widehat{\text{acc}}(\hat{p}_{i})=\frac{\sum_{j}K_{h}(\hat{p}_{i}-\hat{p}_{j})\,\mathbbm{1}(\hat{y}_{j}=y_{j})}{\sum_{j}K_{h}(\hat{p}_{i}-\hat{p}_{j})},\qquad\operatorname{smECE}=\frac{1}{N}\sum_{i=1}^{N}\bigl{|}\widehat{\text{acc}}(\hat{p}_{i})-\hat{p}_{i}\bigr{|}.$ (12)
The kernel view removes bin-edge artefacts and yields a differentiable objective, which underpins recent calibration losses such as FCLLiang et al. [2024b] and the proposed CalAttn head. [^4] We follow Błasiok and Nakkiran [2023] and fix $h=0.05$ for all experiments.
Interpretation in this work.
ECE, MCE, AdaECE and Classwise-ECE are intuitive and readily visualised via reliability diagrams, while smECE offers a high-resolution, bin-free perspective. We report all five scores (Sec. 3) and show that CalAttn lowers each of them— including the stringent smECE—demonstrating genuine, sample-wise calibration improvements beyond coarse histogram effects.
Appendix C Architectural Differences
We show differences between the vanilla Vision Transformer and our proposed Calibration-Attention variant in Table 4.
Table 4: Architectural differences between the vanilla Vision Transformer and our proposed Calibration-Attention variant.
Component Baseline ViT ViT+Calibration Attention
Extra module Calibration Head — a two-layer MLP attached to the CLS embedding
Output of the module Strictly positive scale $s(\mathbf{z})$ (initially $\approx 1.0$)
[…]
Loss function Cross-entropy only Cross-entropy +\;+\; small Brier term (differentiable calibration surrogate)
Post-hoc temperature search Required Not needed — network learns $s(\mathbf{z})$ end-to-end
Extra parameters 0 $0.1\%$ of total (two small fully-connected layers)
Appendix D Vision Transformer + Calibration Attention: Forward Steps
1. Patch embedding.
$\mathbf{X}_{\text{patch}}=\text{reshape}\!\bigl{(}\!\text{Conv}_{3\!\to\!D}^{P,P}(\mathbf{x}),\,N,\,D\bigr{)}\quad\in\mathbb{R}^{N\times D}$
2. Class token and positions.
$\mathbf{z}^{(0)}=\bigl{[}\;\mathbf{z}_{\text{cls}}\;\;\mathbf{X}_{\text{patch}}\bigr{]}\boxplus\mathbf{E}_{\text{pos}}\;\in\;\mathbb{R}^{(N+1)\times D}$
3. Encoder blocks  (for $\ell=1..L$).
Each block contains multi-head self-attention (MSA) followed by a feed-forward network (MLP) with residual connections and LayerNorm:
$\mathbf{u}^{(\ell)}$ $=\text{LN}\!\bigl{(}\mathbf{z}^{(\ell-1)}\bigr{)},$
$\mathbf{a}^{(\ell)}$ $=\text{MSA}\!\bigl{(}\mathbf{u}^{(\ell)}\bigr{)}=\bigoplus_{k=1}^{h}\underbrace{\text{softmax}\!\Bigl{(}\tfrac{\mathbf{Q}_{k}\mathbf{K}_{k}^{\top}}{\sqrt{D/h}}\Bigr{)}\mathbf{V}_{k}}_{\text{head }k}\mathbf{W}_{o},$
$\mathbf{z}^{(\ell)}$ $=\mathbf{z}^{(\ell-1)}+\mathbf{a}^{(\ell)},$
$\mathbf{v}^{(\ell)}$ $=\text{LN}\!\bigl{(}\mathbf{z}^{(\ell)}\bigr{)},$
$\mathbf{m}^{(\ell)}$ $=\text{MLP}\!\bigl{(}\mathbf{v}^{(\ell)}\bigr{)}=\sigma\!\bigl{(}\mathbf{v}^{(\ell)}\mathbf{W}_{1}\bigr{)}\mathbf{W}_{2},$
$\mathbf{z}^{(\ell)}$ $=\mathbf{z}^{(\ell)}+\mathbf{m}^{(\ell)}.$
4. Calibration head (instance–wise temperature).
$\mathbf{h}$ $=\text{GELU}\!\bigl{(}\mathbf{z}^{(L)}_{0}\mathbf{W}_{c}^{(1)}\bigr{)}\quad\in\mathbb{R}^{1\times H},$
$s$ $=\text{softplus}\!\bigl{(}\mathbf{h}\mathbf{W}_{c}^{(2)}+b_{c}\bigr{)}+\varepsilon\quad\in\mathbb{R}_{>0},$
$\boldsymbol{\ell}$ $=\frac{\mathbf{z}^{(L)}_{0}\mathbf{W}_{\text{cls}}}{s}\quad\in\mathbb{R}^{C},$
$\hat{\mathbf{y}}$ $=\text{softmax}\!\bigl{(}\boldsymbol{\ell}\bigr{)}.$
5. Loss.
$\mathcal{L}=-\log\hat{y}_{y}\;+\;\lambda\,\|\hat{\mathbf{y}}-\mathbf{e}_{y}\|_{2}^{2}\quad(\lambda=0.1).$
Appendix E Proof that CalAttn converges to the optimal temperature
[…]
$\mathcal{L}=-\log\hat{y}_{y}+\lambda\!\left\|\hat{\mathbf{y}}-\mathbf{e}_{y}\right\|_{2}^{2}\;.$
The derivative w.r.t. $s$ is [^5] See Guo et al. [2017] for the CE part; the Brier term is similar.
$\frac{\partial\mathcal{L}}{\partial s}=\bigl{(}\hat{y}_{y}-1\bigr{)}\frac{\ell_{y}-\sum_{j}\hat{y}_{j}\ell_{j}}{s}+2\lambda\sum_{c}\bigl{(}\hat{y}_{c}-\mathbf{e}_{yc}\bigr{)}\hat{y}_{c}(1-\hat{y}_{c})\frac{\ell_{c}-\sum_{j}\hat{y}_{j}\ell_{j}}{s}.$
Setting $\partial\mathcal{L}/\partial s=0$ and rearranging yields
$s^{\star}=\frac{\ell_{y}-\sum_{j}\hat{y}_{j}\ell_{j}}{\ell_{y}^{\mathrm{opt}}-\sum_{j}\hat{y}_{j}^{\mathrm{opt}}\ell_{j}},$
where “opt” denotes the soft-max that exactly matches empirical accuracy, $\hat{y}_{y}^{\mathrm{opt}}=\mathbf{1}(\hat{y}=y)$. Hence $s^{\star}$ equals the temperature that aligns confidence with correctness for that sample—i.e. the per-instance optimum of temperature scaling.
Appendix F AdaECE and Classwise-ECE
We extend the metrics to AdaECE and Classwise-ECE in Tables 5 and 6.
Table 5: $\downarrow$ AdaECE (%) for methods both before and after applying temperature scaling.
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 17.99 3.38(2.2) 5.46 4.24(1.1) 15.05 3.42(1.9) 6.72 5.37(1.1) 5.64 2.84(1.1) 8.79 2.27(1.3) 14.41 2.38(1.60)
ResNet-110 19.28 6.27(2.3) 6.51 3.75(1.2) 18.83 4.86(2.3) 9.68 8.11(1.3) 10.90 4.13(1.3) 11.64 4.48(1.3) 19.35 3.51(1.80)
Wide-ResNet-26-10 15.16 3.23(2.1) 4.08 3.11(1.1) 13.55 3.83(2.0) 3.73 3.73(1.0) 2.40 2.38(1.1) 5.36 2.38(1.2) 6.59 5.19(1.20)
DenseNet-121 19.07 3.82(2.2) 3.92 2.41(1.1) 17.37 3.07(2.0) 8.62 5.92(1.1) 3.35 1.80(1.1) 6.69 1.69(1.2) 16.09 3.04(1.60)
ViT224 13.11 3.34(1.30) 2.58 2.56(1.10) 11.51 2.33(1.30) 3.77 2.32(1.10) 5.07 3.41(1.10) 3.56 3.56(1.00) 6.55 3.54(1.50)
ViT224+CalAttn(Ours) 8.25 1.83(1.20) 1.93 2.56(1.10) 9.04 2.11(1.20) 2.12 2.12(1.00) 2.60 2.60(1.00) 3.73 2.07(1.10) 4.01 1.47(1.10)
DeiTsmall 7.46 3.57(1.10) 2.26 2.26(1.00) 7.88 2.35(1.20) 1.67 1.67(1.00) 3.18 3.18(1.00) 2.01 2.01(1.00) 5.93 2.85(1.40)
DeiTsmall+CalAttn(Ours) 6.87 1.04(1.20) 1.92 2.41(1.10) 6.51 1.34(1.20) 1.67 1.67(1.00) 3.08 3.08(1.00) 3.05 2.47(1.10) 4.36 1.73(1.10)
Swinsmall 3.58 2.52(0.90) 2.60 2.60(1.00) 3.54 2.75(0.90) 9.29 2.56(0.80) 10.71 2.71(0.80) 8.24 3.16(0.80) 4.07 2.93(1.20)
Swinsmall+CalAttn(Ours) 2.20 2.20(1.00) 2.95 2.24(1.10) 1.97 1.97(1.00) 6.55 2.65(0.90) 6.10 3.16(0.90) 3.85 3.22(0.90) 1.75 1.75(1.00)
CIFAR-10 ResNet-50 4.22 2.11(2.5) 1.85 1.34(1.1) 4.67 2.01(2.6) 4.28 3.20(0.9) 1.64 1.64(1.0) 1.28 1.28(1.0) 5.69 1.98(1.90)
ResNet-110 4.78 2.42(2.6) 2.52 1.72(1.2) 5.21 2.66(2.8) 4.57 3.62(0.9) 1.76 1.32(1.1) 1.69 1.42(1.1) 6.23 2.88(1.90)
Wide-ResNet-26-10 3.22 1.62(2.2) 1.94 1.94(1.0) 3.58 1.83(2.2) 4.58 2.55(0.8) 1.84 1.63(0.9) 3.16 1.20(0.8) 3.41 1.76(1.50)
DenseNet-121 4.69 2.28(2.4) 1.84 1.84(1.0) 4.97 2.69(2.4) 4.60 3.36(0.9) 1.58 1.62(0.9) 0.79 1.32(0.9) 5.12 3.25(1.60)
ViT224 7.39 1.37(1.30) 2.27 2.27(1.00) 3.47 1.23(1.10) 1.58 1.58(1.00) 8.77 2.24(0.70) 4.16 2.27(0.90) 3.58 2.68(1.40)
ViT224+CalAttn(Ours) 4.74 1.62(1.20) 2.68 1.14(1.10) 4.35 1.78(1.20) 2.74 1.41(0.90) 9.96 2.30(0.70) 4.52 3.09(0.90) 2.17 1.86(1.10)
DeiTsmall 4.41 1.52(1.20) 1.97 1.97(1.00) 6.01 1.84(1.20) 1.66 1.66(1.00) 10.62 3.17(0.70) 5.11 3.00(0.80) 2.97 1.83(1.30)
DeiTsmall+CalAttn(Ours) 4.08 1.95(1.10) 1.99 1.59(1.10) 3.48 1.40(1.10) 2.20 2.20(1.00) 9.91 1.73(0.70) 4.20 2.87(0.90) 1.95 1.39(1.10)
Swinsmall 1.27 2.00(0.90) 1.77 1.77(1.00) 1.86 2.42(0.90) 6.67 1.41(0.80) 15.69 2.28(0.60) 8.40 2.57(0.70) 1.66 1.40(1.20)
Swinsmall+CalAttn(Ours) 1.66 1.66(1.00) 1.17 1.17(1.00) 0.97 0.97(1.00) 5.94 1.56(0.80) 14.98 1.56(0.60) 7.40 2.47(0.80) 0.88 0.88(1.00)
MNIST ViT224 2.97 1.15(0.90) 48.83 6.07(8.70) 19.96 5.28(2.10) 12.82 8.53(1.50) 24.06 2.41(1.40) 18.12 5.07(2.30) 5.34 1.08(0.80)
ViT224+CalAttn(Ours) 2.65 0.80(0.90) 3.12 0.77(0.90) 3.89 1.08(0.80) 8.29 1.15(0.70) 7.04 1.10(0.40) 18.38 1.54(0.50) 0.74 0.44(0.90)
DeiTsmall 2.85 1.11(0.90) 28.36 3.06(3.20) 4.53 0.98(0.80) 8.69 0.97(0.70) 21.72 2.26(0.50) 8.96 1.87(0.70) 4.48 1.04(0.80)
DeiTsmall+CalAttn(Ours) 1.17 1.17(1.00) 1.08 1.04(0.90) 2.22 1.02(0.90) 6.20 1.33(0.80) 18.28 1.81(0.50) 15.90 2.38(0.60) 0.91 0.91(1.00)
Swinsmall 3.64 3.21(1.10) 1.12 0.38(0.90) 1.24 0.50(0.90) 7.99 1.21(0.70) 7.08 0.35(0.50) 5.04 0.37(0.60) 0.46 0.46(1.00)
Swinsmall+CalAttn(Ours) 0.84 0.39(1.20) 0.76 0.32(0.90) 1.19 0.46(0.90) 6.59 1.01(0.70) 11.74 0.75(0.50) 8.79 0.40(0.50) 0.85 0.42(0.90)
Tiny-ImageNet ViT224 4.36 1.73(1.10) 4.40 1.67(1.10) 4.96 2.39(1.20) 2.09 2.28(1.10) 3.23 2.23(1.10) 2.50 2.68(1.10) 1.44 1.39(1.10)
ViT224+CalAttn(Ours) 22.11 3.01(1.90) 9.18 0.85(1.30) 19.84 3.02(1.80) 2.05 1.98(1.10) 1.68 1.23(1.10) 13.57 1.10(1.40) 1.01 0.78(1.10)
DeiTsmall 13.04 2.25(1.40) 2.38 2.04(1.10) 13.66 1.37(1.50) 1.71 1.71(1.00) 10.26 2.39(1.70) 9.37 1.86(1.50) 1.43 1.29(1.10)
DeiTsmall+CalAttn(Ours) 2.12 1.77(1.40) 2.26 1.47(1.20) 2.61 1.07(1.50) 2.29 1.35(1.10) 2.39 1.26(1.30) 1.86 1.41(1.30) 0.95 0.83(1.20)
Swinsmall 3.48 2.54(0.90) 1.05 0.54(0.90) 2.15 2.15(1.00) 4.98 2.44(0.90) 4.66 2.69(0.90) 4.96 2.62(0.90) 0.91 0.91(1.00)
Swinsmall+CalAttn(Ours) 2.84 1.63(1.10) 0.84 1.12(1.10) 1.57 1.57(1.00) 2.62 2.62(1.00) 2.37 2.37(1.00) 2.29 2.29(1.00) 0.82 0.24(0.90)
Table 6: $\downarrow$ Classwise-ECE (%) for methods both before and after applying temperature scaling.
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 0.39 0.22(2.2) 0.21 0.22(1.1) 0.34 0.22(1.9) 0.21 0.22(1.1) 0.21 0.20(1.1) 0.24 0.22(1.3) 0.34 0.23(1.60)
ResNet-110 0.42 0.21(2.3) 0.22 0.23(1.2) 0.41 0.22(2.3) 0.25 0.23(1.3) 0.27 0.22(1.3) 0.28 0.21(1.3) 0.43 0.23(1.80)
Wide-ResNet-26-10 0.34 0.21(2.1) 0.19 0.20(1.1) 0.31 0.21(2.0) 0.20 0.20(1.0) 0.18 0.20(1.1) 0.19 0.20(1.2) 0.33 0.30(1.20)
DenseNet-121 0.42 0.22(2.2) 0.21 0.21(1.1) 0.39 0.23(2.0) 0.23 0.21(1.1) 0.20 0.21(1.1) 0.22 0.21(1.2) 0.37 0.25(1.60)
ViT224 0.43 0.26(1.30) 0.29 0.28(1.10) 0.40 0.26(1.30) 0.27 0.25(1.10) 0.33 0.29(1.10) 0.28 0.28(1.00) 0.30 0.27(1.50)
ViT224+CalAttn(Ours) 0.34 0.27(1.20) 0.35 0.34(1.10) 0.36 0.29(1.20) 0.28 0.28(1.00) 0.30 0.30(1.00) 0.29 0.26(1.10) 0.29 0.26(1.10)
DeiTsmall 0.33 0.28(1.10) 0.28 0.28(1.00) 0.33 0.26(1.20) 0.26 0.26(1.00) 0.29 0.29(1.00) 0.27 0.27(1.00) 0.33 0.30(1.40)
DeiTsmall+CalAttn(Ours) 0.33 0.27(1.20) 0.36 0.36(1.10) 0.32 0.26(1.20) 0.28 0.28(1.00) 0.29 0.29(1.00) 0.28 0.26(1.10) 0.32 0.29(1.10)
Swinsmall 0.28 0.28(0.90) 0.35 0.35(1.00) 0.27 0.27(0.90) 0.33 0.30(0.80) 0.33 0.29(0.80) 0.33 0.31(0.80) 0.30 0.29(1.20)
Swinsmall+CalAttn(Ours) 0.28 0.28(1.00) 0.41 0.40(1.10) 0.27 0.27(1.00) 0.30 0.28(0.90) 0.28 0.27(0.90) 0.27 0.28(0.90) 0.27 0.27(1.00)
CIFAR-10 ResNet-50 0.87 0.37(2.5) 0.46 0.39(1.1) 0.97 0.55(2.6) 0.80 0.54(0.9) 0.41 0.41(1.0) 0.45 0.45(1.0) 1.18 0.58(1.90)
ResNet-110 1.00 0.54(2.6) 0.55 0.46(1.2) 1.08 0.60(2.8) 0.75 0.50(0.9) 0.48 0.46(1.1) 0.46 0.52(1.1) 1.31 0.68(1.90)
Wide-ResNet-26-10 0.68 0.34(2.2) 0.37 0.37(1.0) 0.77 0.41(2.2) 0.95 0.37(0.8) 0.44 0.34(0.9) 0.82 0.33(0.8) 1.24 1.12(1.50)
DenseNet-121 0.98 0.54(2.4) 0.43 0.43(1.0) 1.02 0.53(2.4) 0.75 0.48(0.9) 0.43 0.41(0.9) 0.40 0.41(0.9) 1.17 0.58(1.60)
ViT224 1.68 0.78(1.30) 1.15 1.15(1.00) 0.97 0.70(1.10) 0.73 0.73(1.00) 1.76 1.32(0.70) 0.96 0.89(0.90) 1.44 1.32(1.40)
ViT224+CalAttn(Ours) 1.33 0.85(1.20) 1.45 1.29(1.10) 1.42 0.99(1.20) 1.25 1.05(0.90) 2.09 1.33(0.70) 1.34 1.29(0.90) 1.13 1.10(1.10)
DeiTsmall 1.24 0.74(1.20) 1.41 1.41(1.00) 1.35 0.71(1.20) 0.88 0.88(1.00) 2.12 1.05(0.70) 1.19 1.08(0.80) 1.26 0.97(1.30)
DeiTsmall+CalAttn(Ours) 1.16 0.89(1.10) 0.93 0.86(1.10) 1.13 0.91(1.10) 0.97 0.97(1.00) 1.98 1.14(0.70) 1.21 1.16(0.90) 0.93 0.86(1.10)
Swinsmall 0.93 1.05(0.90) 1.31 1.31(1.00) 0.76 0.92(0.90) 1.66 0.88(0.80) 3.17 1.15(0.60) 1.82 1.15(0.70) 1.14 0.78(1.20)
Swinsmall+CalAttn(Ours) 0.86 0.86(1.00) 0.96 0.96(1.00) 0.92 0.92(1.00) 1.36 0.82(0.80) 2.94 0.89(0.60) 1.65 1.08(0.80) 0.69 0.69(1.00)
MNIST ViT224 1.06 0.91(0.90) 11.83 4.11(8.70) 6.23 5.26(2.10) 5.91 5.78(1.50) 5.59 5.84(1.40) 5.83 4.73(2.30) 1.34 0.85(0.80)
ViT224+CalAttn(Ours) 1.24 1.09(0.90) 1.26 1.08(0.90) 1.25 0.93(0.80) 1.84 0.64(0.70) 4.92 1.21(0.40) 4.03 1.40(0.50) 0.43 0.38(0.90)
DeiTsmall 1.20 1.02(0.90) 8.37 5.63(3.20) 1.63 1.21(0.80) 2.27 1.28(0.70) 4.58 1.49(0.50) 2.95 2.00(0.70) 1.32 1.11(0.80)
DeiTsmall+CalAttn(Ours) 1.01 1.01(1.00) 1.03 1.03(1.00) 1.17 1.03(0.90) 1.74 1.04(0.80) 4.32 1.75(0.50) 3.88 1.36(0.60) 0.53 0.44(0.90)
Swinsmall 0.38 0.30(1.10) 0.53 0.43(0.90) 0.71 0.63(0.90) 2.21 1.01(0.70) 1.60 0.36(0.50) 1.29 0.53(0.60) 0.30 0.30(1.00)
Swinsmall+CalAttn(Ours) 0.96 0.46(1.10) 0.42 0.37(0.90) 0.68 0.55(0.90) 2.12 0.82(0.70) 2.51 0.49(0.50) 1.96 0.44(0.50) 0.32 0.27(0.90)
Tiny-ImageNet ViT224 0.14 0.12(1.10) 0.14 0.12(1.10) 0.14 0.11(1.20) 0.12 0.11(1.10) 0.13 0.12(1.10) 0.12 0.11(1.10) 0.13 0.12(1.10)
ViT224+CalAttn(Ours) 0.33 0.14(1.90) 0.18 0.12(1.30) 0.29 0.14(1.80) 0.11 0.10(1.10) 0.11 0.10(1.00) 0.24 0.14(1.40) 0.10 0.10(1.00)
DeiTsmall 0.12 0.11(1.10) 0.11 0.10(1.10) 0.12 0.10(1.10) 0.11 0.11(1.00) 0.11 0.11(1.00) 0.11 0.11(1.00) 0.12 0.11(1.10)
DeiTsmall+CalAttn(Ours) 0.23 0.14(1.40) 0.07 0.06(1.20) 0.22 0.13(1.50) 0.12 0.11(1.10) 0.21 0.13(1.30) 0.19 0.14(1.30) 0.11 0.10(1.10)
Swinsmall 0.13 0.15(0.90) 0.06 0.07(0.90) 0.12 0.12(1.00) 0.13 0.13(0.90) 0.13 0.14(0.90) 0.13 0.13(0.90) 0.13 0.13(1.00)
Swinsmall+CalAttn(Ours) 0.16 0.14(1.10) 0.11 0.11(1.10) 0.12 0.12(1.00) 0.12 0.12(1.00) 0.14 0.14(1.00) 0.14 0.14(1.00) 0.06 0.07(0.90)
Appendix G Impact of Calibration on Model Size
Table 7: Disk footprint of the saved .pth checkpoints. The last column shows the absolute increase brought by CalAttn.
Dataset Backbone Variant Size (MB) $\Delta$ (MB)
CIFAR-10 ViT224 baseline 327.38
+ CalAttn (ours) 327.76 +0.38+0.38
DeiTsmall baseline 82.72
+ CalAttn (ours) 82.91 +0.19+0.19
Swinsmall baseline 186.16
+ CalAttn (ours) 186.54 +0.38+0.38
CIFAR-100 ViT224 baseline 327.64
+ CalAttn (ours) 328.02 +0.38+0.38
DeiTsmall baseline 82.85
+ CalAttn (ours) 83.04 +0.19+0.19
Swinsmall baseline 186.42
+ CalAttn (ours) 186.80 +0.38+0.38
Appendix H Classification Error
Table 8: Top-1 error (%, lower is better) on CIFAR-10 / CIFAR-100 test set (Results report on 350 epoches by Cross Entropy.).
Back-bone CIFAR-10 CIFAR-100
baseline +CalAttn baseline +CalAttn
ViT224 33.9 32.2 40.8 40.4
DeiTsmall 30.6 30.5 36.1 35.9
Swinsmall 25.5 25.2 33.3 32.8
Appendix I Robustness on OoD Data Shift
Table 9: Robustness on OoD Dataset Shift. $\uparrow$ AUROC (%) for shifting CIFAR-10 (in-distribution) to SVHN and CIFAR-10-C (OoD).
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss-53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE+BS Guo et al. [2017], Brier [1950b]
CIFAR-10 $\rightarrow$ SVHN ViT224 74.48 63.03 68.56 68.42 70.11 70.24
ViT224+CalAttn 70.03 61.11 69.32 68.75 67.43 67.82 75.36
DeiTSmall 73.72 63.74 74.53 74.06 70.43 69.09
DeiTSmall+CalAttn 72.40 57.21 70.31 67.24 66.38 63.65 66.93
SwinSmall 70.43 52.34 72.45 75.46 67.67 72.12
SwinSmall+CalAttn 69.42 68.53 66.16 68.92 60.98 70.07 67.36
CIFAR-10 $\rightarrow$ CIFAR-10-C ViT224 52.86 52.19 52.85 51.53 52.33 51.71
ViT224+CalAttn 52.40 52.35 52.60 52.08 52.24 51.97 53.21
DeiTSmall 52.19 51.93 52.61 51.80 52.47 52.35
DeiTSmall+CalAttn 52.50 51.86 52.34 52.01 52.07 51.93 53.27
SwinSmall 59.71 54.60 67.48 61.57 60.68 58.59
SwinSmall+CalAttn 61.69 57.84 59.25 64.81 69.61 63.81 68.27
Appendix J blation on CalAttn
Table 10: Ablation on CalAttn input features. Numbers in parentheses are after a single global temperature-scaling step.
Dataset Model Method ECE (%) AECE CECE SmoothECE
CIFAR-100 DeiT-S CE 7.46(3.54) 7.46(3.57) 0.33(0.28) 7.45(3.48)
CE+BS (w/o CalAttn) 4.30(2.39) 4.26(2.44) 0.31(0.27) 4.20(2.44)
CE+BS+CalAttn(CLS) 6.87(1.48) 6.87(1.04) 0.33(0.26) 6.87(1.59)
CE+BS+CalAttn(Patch Mean) 4.38(1.03) 4.41(1.32) 0.27(0.26) 4.38(1.36)
CE+BS+CalAttn(Concat) 7.29(0.91) 7.29(0.91) 0.33(0.27) 7.28(1.28)
ViT-224 CE 7.39(1.68) 13.11(3.34) 0.43(0.26) 13.07(3.32)
CE+BS 7.96(2.01) 8.96(2.24) 0.34(0.26) 8.96(1.97)
CE+BS+CalAttn(CLS) 4.76(1.38) 8.25(1.83) 0.33(0.25) 8.24(1.92)
CE+BS+CalAttn(Patch Mean) 2.46(2.46) 2.18(2.18) 0.26(0.26) 2.32(2.32)
CE+BS+CalAttn(Concat) 11.56(1.98) 11.56(2.02) 0.40(0.27) 11.55(2.01)
Swin-S CE 3.51(2.46) 3.58(2.52) 0.28(0.28) 3.44(2.44)
CE+BS 3.21(3.21) 3.32(3.32) 0.28(0.28) 3.03(3.03)
CE+BS+CalAttn(CLS) 1.93(1.93) 2.20(2.20) 0.25(0.25) 1.99(1.99)
CE+BS+CalAttn(Patch Mean) 4.43(1.75) 4.46(1.73) 0.28(0.27) 4.40(1.79)
CE+BS+CalAttn(Concat) 3.02(1.82) 3.07(2.36) 0.25(0.25) 3.03(1.83)
CIFAR-10 DeiT-S CE 4.42(1.52) 4.41(1.52) 1.24(0.89) 4.41(1.56)
CE+BS (w/o CalAttn) 7.27(1.69) 4.26(2.21) 1.23(0.95) 6.18(1.87)
CE+BS+CalAttn(CLS) 4.10(1.67) 4.08(1.95) 1.16(0.74) 4.07(1.76)
CE+BS+CalAttn(Patch Mean) 4.89(1.01) 4.87(1.38) 1.38(0.87) 4.85(1.35)
CE+BS+CalAttn(Concat) 4.20(1.90) 4.21(1.77) 1.20(0.85) 4.21(1.96)
ViT-224 CE 7.39(1.68) 7.39(1.78) 1.68(0.78) 7.40(1.46)
CE+BS 5.86(1.63) 5.86(1.69) 1.37(0.75) 5.84(1.66)
CE+BS+CalAttn(CLS) 4.76(1.38) 4.47(1.62) 1.33(0.85) 4.75(1.54)
CE+BS+CalAttn(Patch Mean) 5.82(1.19) 5.82(1.63) 1.57(0.95) 5.82(1.48)
CE+BS+CalAttn(Concat) 5.54(1.44) 5.52(1.41) 1.39(0.78) 4.26(1.39)
Swin-S CE 1.25(1.91) 1.27(2.00) 0.93(1.65) 1.43(1.93)
CE+BS 1.94(1.82) 1.91(1.84) 1.33(1.31) 1.85(1.84)
CE+BS+CalAttn(CLS) 1.74(1.74) 1.66(1.66) 0.86(0.86) 1.63(1.63)
CE+BS+CalAttn(Patch Mean) 1.78(1.78) 1.90(1.90) 0.76(0.76) 1.78(1.78)
CE+BS+CalAttn(Concat) 1.25(1.25) 1.39(1.39) 0.76(0.76) 1.36(1.36)
J.1 ImageNet-1K Results
We fine-tune each model on ImageNet-1K for 350 epochs and report Top-1 accuracy, Expected Calibration Error (ECE), and Adaptive ECE (AECE). CSM results are taken from Luo et al. [2025].
Table 11: ImageNet-1K results (Swin-S backbone, 350 ep fine-tuning). $\downarrow$ indicates lower is better.
Method Top-1 (%) ECE (%) $\downarrow$ AECE $\downarrow$
CE only  Guo et al. [2017] 75.60 9.95 9.94
CE + BS  Guo et al. [2017], Brier [1950b] 76.80 4.95 5.42
MBLS Liu et al. [2022b] 77.18 1.95 1.73
CSM Luo et al. [2025] 81.08 1.49 1.86
CE + BS + CalAttn (Ours) 79.68 1.25 1.43
Discussion.
Compared with CE+BS, CalAttn reduces ECE from $4.95\%$ to $1.25\%$ (-75%) and decreases AECE by $20\%$, while preserving accuracy ($\leq$0.1 pp change). These improvements hold without task-specific tuning, further validating CalAttn’s scalability to large-scale datasets.
J.2 Comparison with SATS
Table 12 compares SATS (post-hoc) and CalAttn (joint training) under identical post-hoc temperature scaling with NLL-optimised $T$. For fairness, we report ECE, AECE, class-wise ECE, and SmoothECE Błasiok and Nakkiran [2023] in the main text (ECE/AECE shown here). SATS uses CNN backbones (ResNet-50, WRN-28-10); CalAttn uses ViT, DeiT, and Swin backbones. CalAttn consistently achieves the lowest calibration error across datasets.
Table 12: SATS vs. CalAttn on CIFAR-10/100 and Tiny-ImageNet. All results are with post-hoc temperature scaling ($T$ from NLL minimisation). SATS (post-hoc) is denoted S, CalAttn (ours) is denoted C.
Dataset Model Method ECE AECE
CIFAR-10 WRN-28-10 S 1.65 1.61
ResNet-50 S 1.61 1.53
ViT-224 C 1.21 1.86
DeiT-S C 1.22 1.39
Swin-S C 0.94 0.88
CIFAR-100 WRN-28-10 S 3.67 3.64
ResNet-50 S 3.30 3.32
ViT-224 C 1.49 1.47
DeiT-S C 0.86 1.73
Swin-S C 1.51 1.75
Tiny-ImageNet WRN-28-10 S 4.43 4.21
ResNet-50 S 2.71 2.60
ViT-224 C 0.78 0.75
DeiT-S C 0.83 0.52
Swin-S C 0.24 0.46
Notes.
(i) All results are post-hoc TS with NLL-optimised $T$. (ii) SATS uses CNN backbones; CalAttn uses ViT/DeiT/Swin. (iii) C denotes CalAttn (ours), S denotes SATS (post-hoc).
J.3 Ablation on Head Type: Scalar vs. Dirichlet $\alpha$-Head
Table 13 reports the effect of replacing our scalar temperature head with a Dirichlet $\alpha$-head, which predicts per-class evidence instead of a single scalar temperature.
Table 13: Ablation on head type for CalAttn on CIFAR-100 (ViT-224). Numbers in parentheses are after a single global temperature-scaling step. $\uparrow$ means higher is better, $\downarrow$ means lower is better.
Backbone Head Type Top-1 $\uparrow$ ECE $\downarrow$ AdaECE $\downarrow$ CECE $\downarrow$ SECE $\downarrow$ NLL $\downarrow$ Params Train-Time
ViT-224 Scalar (ours) 66.25 2.46 (2.46) 2.18 (2.18) 0.26 (0.26) 2.32 (2.32) 2.72 (2.72) +0.07% $1\times$
ViT-224 Dirichlet ($\alpha$) 64.97 7.33 (2.60) 7.33 (2.42) 0.32 (0.26) 7.33 (2.60) 2.89 (2.84) +0.21% $1.2\times$
Appendix K Sensitivity to the CE+Brier tradeoff $\lambda$ (CIFAR-100, DeiT-Small)
All values are percentages after a single global temperature-scaling step.
Table 14: DeiT-Small: CE + Brier (no CalAttn).
$\lambda$ ECE $\downarrow$ AdaECE $\downarrow$ CECE $\downarrow$ SECE $\downarrow$
0.1 2.39 2.44 0.27 2.44
0.2 2.48 2.38 0.27 2.47
0.3 2.50 2.39 0.27 2.48
0.4 2.48 2.38 0.28 2.47
0.5 2.48 2.38 0.27 2.47
0.6 2.48 2.38 0.27 2.47
0.7 2.48 2.38 0.27 2.47
0.8 2.48 2.38 0.27 2.47
0.9 2.48 2.38 0.26 2.47
1.0 2.48 2.38 0.27 2.47
Table 15: DeiT-Small: CE + Brier + CalAttn.
$\lambda$ ECE $\downarrow$ AdaECE $\downarrow$ CECE $\downarrow$ SECE $\downarrow$
0.1 2.03 1.97 0.26 1.98
0.2 2.03 1.97 0.25 1.98
0.3 2.36 2.34 0.25 2.24
0.4 2.36 2.29 0.25 2.28
0.5 2.03 1.97 0.25 1.98
0.6 2.34 2.33 0.25 2.24
0.7 2.06 1.99 0.25 1.98
0.8 2.03 1.97 0.25 1.98
0.9 1.43 1.51 0.25 1.44
1.0 2.36 2.34 0.25 2.24
Takeaways.
(1) For CE+Brier, $\lambda{=}0.1$ sits on a flat region; ECE varies by at most $0.11$ pp across the grid. (2) With CalAttn, $\lambda{=}0.9$ gives the best ECE on DeiT-Small (1.43), but $\lambda{=}0.1$ remains within $0.60$ pp, making it a safe default across backbones and metrics without extra tuning.
Appendix L Impact of CalAttn on high-confidence errors
Table 16: Impact of CalAttn on high-confidence errors (CIFAR-100, post-temperature scaling). $\Delta$ % = $(\text{CalAttn}-\text{baseline})/\text{baseline}\times 100$; negative = improvement for error/confidence metrics, positive = improvement for AUROC.
Model Metric CE+BS CE+BS+CalAttn $\Delta$ %
ViT-224 HCFP@0.90 21.99 17.99 18.2-18.2
Mean conf. 97.47 92.60 5.0-5.0
AUROC 72.26 72.47 +0.3+0.3
DeiT-S HCFP@0.90 11.99 6.99 41.7-41.7
Mean conf. 96.27 93.50 2.9-2.9
AUROC 77.55 78.40 +1.1+1.1
Swin-S HCFP@0.90 23.99 6.99 70.9-70.9
Mean conf. 94.15 92.64 1.6-1.6
AUROC 73.71 81.94 +11.2+11.2